# Boulevard du Triomphe, B-1050 Brussels, Belgium

### v1

@@ 3.1 Upper Limits
Figure 6: Constraints on the characteristic strain $h_{c}$ from different experiments. Blue and green solid lines represent our bounds from an induced photon flux in the magnetosphere of the galactic NSs, for NS models with a constant and decaying $\vec{B}$, respectively. We also show limits from OSQAR, ALP and CAST Ejlli:2019bqj as well as the one from ARCADE (the upper and lower lines corresponding to uncertainty on the cosmic magnetic field) Domcke:2020yzq and Big Bang Nucleosynthesis Cyburt:2015mya (black dashed-dotted lines). The red and orange dashed curves represent the limits from conversion in the geomagnetic field and galactic magnetic field respectively Ito:2023nkq .
The limits found with this method are reported as blue and green solid lines in Fig. 6, for a constant and a decaying magnetic field, respectively. For comparison, we also show the current limits derived from axion experiments: OSQAR, ALP and CAST Ejlli:2019bqj as well as those derived using ARCADE Domcke:2020yzq (see Refs.Domcke:2023qle ; Aggarwal:2020olq for a summary of those constraints). The orange and red dashed curves show the upper limits derived when comparing the induced photon-flux produced in the geomagnetic field and galactic magnetic field Ito:2023nkq to the same telescope observations used in this work. Interestingly, we find our limits to be the best in the frequency range $(10^{8}-10^{9})\,\mathrm{Hz}$ and comparable to existing bounds around $(10^{12}-10^{15})\,\mathrm{Hz}$. In contrast with bounds from geomagnetic conversion, our results predict a flux which peaks in the Galactic Center, as shown in Fig. 5. Observations of the galactic flux angular distribution or focused on the Galactic Center could improve our bounds with respect to others. At larger frequencies, the induced photon flux from galactic NSs is systematically smaller than the one from the geomagnetic field and galactic magnetic field. For this reason our limits are less competitive at large frequencies.
@@
Finally, the black dashed-dotted line shows the upper limit on the amount of gravitational waves produced before Big Bang Nucleosynthesis (BBN), which would contribute to the number of relativistic degrees of freedom Cyburt:2015mya . Since this limit does not apply to post-BBN or monochromatic sources, current bounds are important even if they fall short compared to BBN constraints. Still, these bounds correspond to GWs with $\Omega_{\mathrm{gw}}\gg 1$. Even if exotic, such abundances could be produced according to models of primordial black holes mergers with mass $\sim(10^{-7}-10^{-9})M_{\odot}$ Domcke:2020yzq ; Maggiore:2007book ; Franciolini:2022htd .
@@ Acknowledgements
We thank Valerie Domcke and Jordi Salvadó for useful comments on the manuscript and Asuka Ito, Kazunori Kohri, Kazunori Nakayama for useful discussion. TB acknowledges financial support from the Spanish grants PID2022-136224NB-C21, PID2019-108122GBC32, PID2019-105614GB-C21, and from the State Agency for Research of the Spanish Ministry of Science and Innovation through the “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences (CEX2019-000918-M).
@@ Appendix A Neutron Star Distribution
In order to accurately predict the total induced photon flux on Earth, a realistic distribution for the properties of the NSs must be used (surface magnetic field $B$  [^4] Later on, $B$ will always refer to the surface magnetic field. and spin period $T$).
Hopefully, numerous NS observations have been performed (more than 2000 rotation-powered NSs Manchester:2001fp ; Edwards:2001de ). Based on those, recent efforts have been made in developping NS models (initial magnetic field and period distributions as well as their time evolution) capable of reproducing the observed data. In this appendix, we summarized the ideas of two distinct models that are used in the main text Faucher-Giguere:2005dxp ; Popov:2009jn .
In the first (Model 1, referred in the main text to ”constant magnetic field”), the NS magnetic field is assumed constant and therefore fixed to its initial birth value. On the other hand, the spin period is increasing in time as the NS loses energy by dipole radiation and plasma effects (see Refs.Spitkovsky:2006np ; Philippov:2013aha ; Johnston:2017wgm ). It is found that the period evolves as Safdi:2018oeu
@@
Eventually, as the period increases, the NS spin might be low enough to prevent a coherent radio pulsed emission. At this stage, the NS becomes undetectable (further on referred as ”dead” or ”inactive” NSs, although it is important to stress that it does not imply that it loses its magnetic field) Zhang:2000rd . This transition is expected to happen when the ratio $B/T^{2}$ is smaller than Bhattacharya:1992vf
@@
Post-death, NSs are expected to lose energy only through dipole radiation Safdi:2018oeu , this translate into a modified period evolution,
@@
Model 1
$\sigma_{B}=0.55,\,\sigma_{T}=0.15$
$\mu_{B}=12.95,\,\mu_{T}=0.3$
Model 2
$\sigma_{B}=0.6,\,\sigma_{T}=0.1$
$\mu_{B}=13.25,\,\mu_{T}=0.25$
@@
In the second model (Model 2, referred in the main text to ”decaying magnetic field”), the spin period time evolution is considered to be the same as in Model 1. Nonetheless, the magnetic field is here considered decaying in time. This is the essential difference between these two models. For the magnetic field decay, we follow Ref.Safdi:2018oeu and describe the time evolution of the surface magnetic field as
@@
Furthermore, from the same simulations, the today’s magnetic field and period distributions can be extracted. In both models, it has been found that only $\sim 0.4\%$ of the NSs are still active Safdi:2018oeu . In Fig.7, we show the today’s distributions for both the surface magnetic field and the spin period in the two models. In particular, we show separately the distributions (normalized) for the active and dead NSs. It can be appreciated that, in both models, the dead NSs have larger periods. This is directly related to the fact that dead NSs have more time to evolve their spin period to larger values. On the rigth pannel, the magnetic field distributions for the active and dead NSs in the Model 1 are relatively similar with slighly larger values for the dead NSs. This is due to the fact that for larger magnetic fields, $\tau_{0}$ gets smaller making the spin period descreases faster. Interestingly, the magnetic field distributions for the Model 2 are quite different. Recalling that in this model, the magnetic field is assumed to decay in time, the dead NSs, assumed to be older than the active ones, have more time to decay and for this reason their surface magnetic field is significantly reduced compared to the active population (assumed younger).
@@
The total number of NSs can be estimated from the chemical composition of the Galaxy. In particular, its iron content provides the information that about $10^{9}$ NS have been produced Ofek:2009wt . Similar estimation can be found from the rate of core-collapse supernovae observed in our current days and assuming that this rate has been constant over the lifetime of the galaxy Sartore:2009wn .
In line with this, for both models the NSs have been splitted into a bulge and a disk population with respectively $N_{\rm bulge}=4.8\times 10^{8}$ and $N_{\rm disk}=3.2\times 10^{8}$ NSs Ofek:2009wt ; Sartore:2009wn . More precisely, the disk distribution (in a frame with the origin at the galactic center) has been extracted from millisecond pulsar observations Lorimer:2006qs ; Bartels:2018xom (see also Faucher-Giguere:2009fex for another parametrisation),
@@
As already mentionned, from this spatial distribution, $0.4\%$ will be active NSs with surface magnetic field and spin period distributed as in Fig.7 (green solid lines) and $99.6\%$ will be inactive NSs with surface magnetic field and spin period distributed as the blue dashed lines in Fig.7. In the main text, the total flux on Earth from graviton-to-photon conversion in the NS magnetospheres has been calculated for both Model 1 and Model 2.

### v2

@@ 3.1 Upper Limits
[…]
The limits found with this method are reported as blue and green solid lines in Fig. 6, for a constant and a decaying magnetic field, respectively. For comparison, we also show the current limits derived from axion experiments: OSQAR, ALP and CAST Ejlli:2019bqj as well as those derived using ARCADE Domcke:2020yzq (see Refs.Domcke:2023qle ; Aggarwal:2020olq for a summary of those constraints). The orange and red dashed curves show the upper limits derived when comparing the induced photon-flux produced in the geomagnetic field and galactic magnetic field Ito:2023nkq to the same telescope observations used in this work. Interestingly, we find our limits to be the best in the frequency range $(10^{8}-10^{9})\,\mathrm{Hz}$ and comparable to the geomagnetic bounds around $(10^{12}-10^{15})\,\mathrm{Hz}$. In contrast with bounds from geomagnetic conversion, our results predict a flux which peaks in the Galactic Center, as shown in Fig. 5. Observations of the galactic flux angular distribution or focused on the Galactic Center could improve our bounds with respect to others. At larger frequencies, the induced photon flux from galactic NSs is systematically smaller than the one from the geomagnetic field and galactic magnetic field. For this reason our limits are less competitive at large frequencies.
@@
Finally, the black dashed-dotted line shows the upper limit on the amount of gravitational waves produced before Big Bang Nucleosynthesis (BBN), which would contribute to the number of relativistic degrees of freedom Cyburt:2015mya . Since this limit does not apply to post-BBN or monochromatic sources, current bounds are important even if they fall short compared to BBN constraints. Still, these bounds correspond to GWs with $d\Omega_{\mathrm{gw}}/d\log f\gg 1$. Such sources would need to be extremely monochromatic in order not to overclose the Universe. Even if they emitted in a more reasonable frequency width, no known cosmological sources can produce HFGWs intense enough to saturate our bounds. The mergers of exotic compact objects or cosmological sources could produce GWs with the right frequency, but with $\Omega_{\mathrm{gw}}<1$ Giudice_2016 ; Domcke:2020yzq ; Maggiore:2007book ; Franciolini:2022htd ; Aggarwal:2020olq . Therefore, further improvement is still required for all state-of-the-art HFGWs analysis to constrain such sources.
With respect to our results, the bounds here presented will be improved by more precise all-sky observations. It might also be possible to profit from the accumulation of neutron stars in the Galactic center direction. However, such backgrounds are harder to model and a lower signal-to-noise ratio would make an improvement challenging.
Again, these kind of analysis would only constrain homogeneous fluxes, either of cosmological or astrophysical (but distributed around all the sky) origin. Single localized events would require a perfect alignment between the NS and the GW source and require a different approach to an all-sky observation. Thus, these are beyond the scope of our work.
@@ Acknowledgements
We thank Valerie Domcke, Jordi Salvadó and Jinsu Kim for useful comments on the manuscript and Nicolas Grimbaum, Asuka Ito, Kazunori Kohri, Kazunori Nakayama for useful discussion. TB acknowledges financial support from the Spanish grants PID2022-136224NB-C21, PID2019-108122GBC32, PID2019-105614GB-C21, and from the State Agency for Research of the Spanish Ministry of Science and Innovation through the “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences (CEX2019-000918-M). The work of F.C. is supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 860881-HIDDeN.
@@ Appendix A Neutron Star Distribution
To accurately predict the total induced photon flux on Earth, a realistic distribution for the properties of the NSs must be used (surface magnetic field $B$  [^4] Later on, $B$ will always refer to the surface magnetic field. and spin period $T$).
Numerous NS observations have been performed (more than 2000 rotation-powered NSs Manchester:2001fp ; Edwards:2001de ). Based on those, recent efforts have been made in developing NS models (initial magnetic field and period distributions as well as their time evolution) capable of reproducing the observed data. In this appendix, we summarized the ideas of two distinct models that are used in the main text Faucher-Giguere:2005dxp ; Popov:2009jn .
In the first (Model 1, referred to in the main text as ”constant magnetic field”), the NS magnetic field is assumed constant and therefore fixed to its initial birth value. On the other hand, the spin period increases in time as the NS loses energy by dipole radiation and plasma effects (see Refs.Spitkovsky:2006np ; Philippov:2013aha ; Johnston:2017wgm ). It is found that the period evolves as Safdi:2018oeu
@@
Eventually, as the period increases, the NS spin might be low enough to prevent a coherent radio-pulsed emission. At this stage, the NS becomes undetectable (further on referred to as ”dead” or ”inactive” NSs, although it is important to stress that it does not imply that it loses its magnetic field) Zhang:2000rd . This transition is expected to happen when the ratio $B/T^{2}$ is smaller than Bhattacharya:1992vf
@@
Post-death, NSs are expected to lose energy only through dipole radiation Safdi:2018oeu , this translates into a modified period evolution,
@@
Model 1 $\sigma_{B}=0.55,\,\sigma_{T}=0.15$ $\mu_{B}=12.95,\,\mu_{T}=0.3$
Model 2 $\sigma_{B}=0.6,\,\sigma_{T}=0.1$ $\mu_{B}=13.25,\,\mu_{T}=0.25$
@@
In the second model (Model 2, referred in the main text as ”decaying magnetic field”), the spin period time evolution is considered to be the same as in Model 1. Nonetheless, the magnetic field is here considered decaying in time. This is the essential difference between these two models. For the magnetic field decay, we follow Ref.Safdi:2018oeu and describe the time evolution of the surface magnetic field as
@@
Furthermore, from the same simulations, the today’s magnetic field and period distributions can be extracted. In both models, it has been found that only $\sim 0.4\%$ of the NSs are still active Safdi:2018oeu . In Fig.7, we show today’s distributions for both the surface magnetic field and the spin period in the two models. In particular, we show separately the distributions (normalized) for the active and dead NSs. It can be appreciated that, in both models, the dead NSs have larger periods. This is directly related to the fact that dead NSs have more time to evolve their spin period to larger values. On the right panel, the magnetic field distributions for the active and dead NSs in Model 1 are relatively similar with slightly larger values for the dead NSs. This is due to the fact that for larger magnetic fields, $\tau_{0}$ gets smaller making the spin period decrease faster. Interestingly, the magnetic field distributions for the Model 2 are quite different. Recalling that in this model, the magnetic field is assumed to decay in time, the dead NSs, assumed to be older than the active ones, have more time to decay and for this reason, their surface magnetic field is significantly reduced compared to the active population (assumed younger).
@@
There exist different methods to determine the number of neutron stars in the Milky Way. On the one hand, supernovae explosions, which lead into to NS, release an important amount of iron. The study of the chemical composition of the galaxy in the Milky Way and its iron content suggests that about $10^{9}$ NSs have been produced Arnett1989 . On the other hand, the disk stellar birth rate has been observed to be approximately constant over the galactic history Rocha-Pinto2000 . This implies a roughly constant NS formation rate as well, which leads to $4\times 10^{8}$ NSs in the disk. The stellar population in the bulge is expected to have formed in a much shorter time scale around 12 Gyrs ago, and its stellar mass function is skewed toward high masses Ballero2007 . The early time population of heavy stars led to a more important NS population in the bulge, around $6\times 10^{8}$ Ofek:2009wt . All in all, this leads again to the estimation of $10^{9}$ NSs in the galaxy. We follow standard choices in the literature and take $N_{\rm bulge}=4.8\times 10^{8}$ and $N_{\rm disk}=3.2\times 10^{8}$ NSs Ofek:2009wt ; Sartore:2009wn . Of course, depending of the total number of NSs, the upper limits derived in this work would change. It is expected that increasing (decreasing) the number of NSs would translate into stronger (weaker) constraints.
Each population follows a different spatial distribution. The disk distribution (in a frame with the origin at the galactic center) has been extracted from millisecond pulsar observations Lorimer:2006qs ; Bartels:2018xom (see also Faucher-Giguere:2009fex for another parametrisation),
@@
As already mentioned, from this spatial distribution, $0.4\%$ will be active NSs with surface magnetic field and spin period distributed as in Fig.7 (green solid lines) and $99.6\%$ will be inactive NSs with surface magnetic field and spin period distributed as the blue dashed lines in Fig.7. In the main text, the total flux on Earth from graviton-to-photon conversion in the NS magnetospheres has been calculated for both Model 1 and Model 2. For both models, we assumed an NS radius of $r_{0}=10$ km REF.